# Learning Geospatial Region Embedding with Heterogeneous Graph

### v1

@@ 1 Introduction
Geospatial regions, as complex systems, witness the intertwining of natural laws and social dynamics, leading to mixed-order interactions among regions and entities. Recently, there has been a growing interest in leveraging deep learning methods to tackle various geospatial tasks [59, 41], such as regional indicator prediction [43, 42, 45, 14], earth monitoring [19, 25, 47], and geopolitics optimization [12, 31, 40]. Global geospatial regions exhibit intricate and multifaceted dynamics. Learning effective geospatial embeddings that capture the inherent characteristics of regions and their intricate mixed-order relationships lays the groundwork for these applications [55, 33, 19].
To learn effective representations for geographical regions, it is essential to capture two key characteristics: intra-region features and inter-region correlations [30, 53, 5]. As illustrated in Figure 1, (a) intra-region features encapsulate a region’s environmental information (e.g., vegetation coverage, and water resources) and societal information (e.g., colleges, or government institutions). (b) Inter-region correlations indicate the relationships between regions, extending beyond simple pairwise adjacency to capture higher-order dependencies. Specifically, we believe that the representation of a region can be influenced by multiple regions collectively, even if they are not spatially adjacent. For instance, regions along the same river basin may exhibit correlated climate and economic patterns despite lacking direct adjacency.
Figure 1: Illustration of region representation from intra-region and inter-region perspective. (a) intra-region features consist of environmental features and societal features. (b) inter-region correlations, e.g., the second-order adjacent correlation between neighboring regions; and the high-order correlation between remote region groups induced by environmental and societal factors.
Although relevant works [35, 42, 21] have sought to learn better geospatial representations, concisely learning an effective and comprehensive geospatial embedding remains challenging. The questions are two-fold. First, an effective representation for intra-region features remains unexplored. Existing studies opt for leveraging globally available satellite imagery to learn geospatial representations that can generate a unified embedding space for regions across the world [35, 42, 20]. These studies often employ vision-based encoders pre-trained on natural images to learn semantics from satellite imagery, but such encoders struggle due to the substantial differences between satellite and natural imagery. Second, comprehensively modeling the complex high-order inter-region relations in geospatial embeddings is challenging. In geo-space, inter-region relationships extend beyond second-order adjacency to encompass ternary and higher-order dependencies between regions and region groups. For instance, a certain region may be influenced by several non-adjacent regions due to shared geographic or socioeconomic functions, despite lacking physical proximity [30, 21]. While multi-view graphs have been explored to model complex geographic and socioeconomic relations, they heavily rely on specific data to construct and are limited to second-order perspectives [59, 16, 17]. Alternatively, knowledge graphs (KGs) [27] built from geographical data, such as POIs, aim to capture complex structures, but their complexity limits scalability and effectiveness in training [27, 17].
To address these challenges, we propose a novel Heterogeneous Graph structure with a learning framework for effective and mixed-order relation-aware Geospatial embedding (named GeoHG). Our framework leverages satellite images and POI information to effectively derive unified region representations encapsulating both intra-region and inter-region information from environmental and societal perspectives. To derive effective intra-region representations that generalize across global regions, we proposed a novel satellite image encoding mechanism. Instead of directly using vanilla vision encoders pre-trained on life images, which struggle with the domain gap, we perform semantic segmentation on the satellite images to effectively distinguish geo-entities like water bodies, vegetation, and man-made structures. Subsequently, we integrate the spatial coordinate with the geo-entities and extra POI information within each region to construct a comprehensive intra-regional feature representation. To efficiently capture high-order inter-region relations, we leverage the powerful mixed-order dependencies from the heterogeneous graph that reflects both environmental and societal aspects of geographical regions. The heterogeneous formulation integrates spatial attributes with explicit region-entity associations, representing the high-dimensional socio-environmental dependencies across regions in a unified framework. Finally, our GeoHG seamlessly integrates the intra-region and inter-region representations through model-agnostic heterogeneous graph structure. Its universal compatibility allows for seamless integration with different models and algorithms for various tasks, ensuring optimal performance across diverse applications with effective geospatial representation.
To summarize, the contributions of this work are concluded as follows:
  • Novel Heterogeneous Graph Structure for Geospatial Embedding. We introduce a model-agnostic heterogeneous graph structure to integrate intra-region features and inter-region correlations in geospace efficiently. As far as we know, it is the first work that integrates comprehensive intra-region information with complex mixed-order inter-region relations in a heterogeneous graph for geospatial representation (Sec 3.3).
  • Efficient and Explicit Intra Region Embedding. We develop an effective approach to construct intra-region feature representation by leveraging entity segments from satellite images to extract environmental features and utilizing POI information to capture societal attribute features (Sec 3.1).
  • Mixed-order Inter-region Relation-Aware Representation. We propose a novel method to construct interpretable heterogeneous graphs that explicitly capture mixed-order relations between regions. Our approach effectively represents the intricate multivariate relationships within the regions’ natural environment. Additionally, it captures the complex social attribute relationships among various types of social entities within geospatial regions (Sec 3.2).
  • Empirical Evidence. Our empirical studies across various downstream tasks in different locations demonstrate the superior performance of GeoHG compared to various baseline models. It consistently outperforms existing methods and maintains substantial gains even in extremely low-data regimes (5% training data). Moreover, GeoHG outperforms conventional geospatial interpolation methods specifically designed for such data-scarce settings (Sec 4.1).

### v2

@@ 1 Introduction
Our living space, as a complex system, reflects the interplay between natural laws and social dynamics. Socioeconomic indicators are designed to quantify the societal and economic dynamics of the space and contribute significantly to various domains such as the environment, management, sustainability, and urban computing (Zou et al., 2025; Xie et al., 2020). In practice, these indicators are typically acquired through field investigations (e.g., population census) or deployed equipment (e.g., air monitoring station), with limited budgets resulting in a number of ground truth samples that is much lower than the global demand (Costa et al., 2012; Ross and Reeve, 2003). In this context, Socioeconomic Indicator Inference aims to estimate the values of uninvestigated regions based on limited observations in order to provide a more uniform and comprehensive quantification of the overall geospace.
Figure 1: Insight for space-aware no-continuous indicator inference.
In the past decade, most socioeconomic indicator inference methods, such as the kriging (Brus and Heuvelink, 2007) method, have been developed using geometric approaches for spatial interpolation (Mitas and Mitasova, 1999). While feasible, these methods depend on key assumptions of spatial continuity, spatial stationarity, and isotropic spatial autocorrelation, which imply that the events of the considered spatial indicators are distributed uniformly across the region, regardless of geographical directions or barriers (Mitas and Mitasova, 1999; Tan and Xu, 2014). However, the actual distribution of these indicators is non-continuous and influenced by multiple dependencies. For instance, two adjacent regions may have different population densities due to variations in land environment (Ross and Reeve, 2003), while regions along the same river basin may exhibit similar GDP patterns despite being separated by distance (Li et al., 2024; Wang et al., 2019). Consequently, the interpolated values can significantly diverge from the real scenario. Additionally, the available samples are typically non-uniformly distributed in space, resulting in poor or even ineffective performance of traditional interpolation methods in sparse areas (Li and Heap, 2011).
Figure 2: Illustration of geospatial features and dependencies for spare-aware indicator inference.
To address this issue, it is important to integrate artificial intelligence to uniformly represent features and non-Euclidean relationships within geospace using geospatial data. This representation can then enhance global indicator inference in non-continuous and non-uniform scenarios (see Fig. 1). While significant research on region representation exists (Liu et al., 2023; Wang et al., 2020; Hao et al., 2025; Zhang et al., 2024; Mai et al., 2024), most studies assume redundant samples and train models on large datasets to predict individual samples. The design of a method that comprehensively represents geospatial data for indicator inference remains unexplored.
Introducing such methods presents two main challenges. First, the data reliance for geospace representation must be robust. Existing region representation methods may utilize various data sources, such as trajectories, street-view imagery, or road networks, which might not always be available. Although some studies leverage globally available satellite imagery to generate general spatial representations, their specified vision encoders may perform poorly when trained with only a few samples. Second, comprehensively representing environmental and societal dependencies in geospatial data is non-trivial. As shown in Figure 2, space-aware inference should jointly consider spatial geometry, environmental, and societal dependencies, which extend beyond basic regional features and second-order adjacency to include high-order dependencies across regions. For instance, a certain region may be influenced jointly by a group of non-adjacent regions due to shared geographic or socioeconomic functions.
To address these challenges, we present GeoHG, the first space-aware socioeconomic indicator inference method that utilizes a heterogeneous graph-based structure to jointly represent geospace from spatial geometry, environmental, and societal perspectives for non-continuous inference. For robustness, we employ only satellite imagery and points of interest (POI) to represent the environmental and societal features of the space, as these are among the most accessible data sources. Furthermore, to overcome the limitations of traditional vision encoders, we directly perform semantic segmentation on the satellite imagery to distinguish geo-entities such as water bodies, vegetation, and buildings, enabling a more general and concise representation. For comprehensive geospace representation, we introduce a novel heterogeneous graph that integrates spatial geometrical relations with encoded environmental and societal information. This graph structure effectively captures the non-Euclidean local relationships within the space. Additionally, our heterogeneous formulation between region-entity associations accounts for the high-dimensional societal and environmental dependencies across regions. By uniformly representing the space, we can conduct inference on the graph for the entire region with consistent accuracy, regardless of the distribution of available samples. Overall, our contribution can be summarized as follows:
  • Space-aware Inference. We introduce a new space-aware method for socioeconomic indicator inference that incorporates a joint representation of spatial geometry, environmental, and societal relations. To the best of our knowledge, this is the first work that integrates graph-based space representation for socioeconomic indicator inference.
  • Heterogeneous Graph for Joint Space Representation. We develop a heterogeneous graph structure to model geospace using widely accessible data sources, capturing high-order societal and environmental dependencies through heterogeneous connections.
  • Empirical Evidence. Our experiments demonstrate that GeoHG outperforms existing interpolation methods and region representation models across various downstream tasks. It achieves an $R^{2}$ score exceeding 0.8 under extreme data scarcity with a masked ratio of 95% and performs effectively on inference with highly unevenly distributed samples.
Figure 3: Overall framework of the proposed GeoHG.
2 Formulation
An investigated geospace $\mathcal{R}$ typically comprises multiple regions denoted as =$\mathcal{R}=\left\{r_{1},r_{2},\cdots,r_{N}\right\}$. These regions exhibit unique environmental and societal characteristics, often reflected through various data sources within them. In our study, we represent $r_{i}\in\mathcal{R}$ through the following aspects:
  • Spatial Geometry represents the spatial location and geometrical relations of $r_{i}$. The spatial location can be represented by its center coordinate $(lon^{i},lat^{i})$, and the geometrical relations can be captured by computing the distance between each pair of regions $r_{i}$.
  • Satellite Imagery captured through ongoing earth observation by satellites, provides an overview of the environmental features of $\mathcal{R}$. Each satellite image of a region $r_{i}$ can be expressed as $S_{i}\in\mathbb{R}^{H\times W\times L}$, where $H\times W$ denotes the image’s size and $L$ represents the number of color channels.
  • Point of Interest refers to a specific location or geographic point of significance to society, which includes campuses, businesses, or any other notable locations. POI can be represented as $P=(x,y,c)$, where $x$ and $y$ are the geographical coordinates, and $c$ is the societal category.
Problem Statement. Given a geospace $\mathcal{R}$ with corresponding satellite images $\mathbf{S}=\left\{S_{1},S_{2},\cdots,S_{N}\right\}$ and a set of POIs $\mathbf{P}_{i}=\left\{P_{1},P_{2},\cdots,P_{M}\right\}$, our goal is to construct a graph $\mathcal{G}$ that captures the spatial and contextual characteristics of the space. Within the space graph $\mathcal{G}$, our approach aims to generate precise estimates of socioeconomic indicators $\mathcal{Y}$ for the unlabeled regions $\mathcal{R}_{mask}\subseteq\mathcal{R}$ from the available regions $C_{\mathcal{R}_{mask}}\mathcal{R}$. This process can be formulated as:
$\left(\hat{\mathcal{Y}}^{r_{1}},\hat{\mathcal{Y}}^{r_{2}},\cdots,\hat{\mathcal{Y}}^{r_{k}}\right)\leftarrow\mathcal{F}_{\theta}\left(\mathcal{G}|\mathcal{Y}^{r_{k+1}},\mathcal{Y}^{r_{k+2}},\cdots,\mathcal{Y}^{r_{N}}\right),$
where $\{r_{1},r_{2},\cdots,r_{k}\}\in\mathcal{R}_{mask}$ and we use $\mathcal{M}=k/N$ as the masked ratio. $\mathcal{F}_{\theta}$ is the inference model.
3 Methodology
As illustrated in Figure 3, our GeoHG framework comprises four major components:
  • Space Feature Representation. We first extract basic space features from satellite imagery and POI data, considering the proportion of geo-entities, POI category distribution, and the regions’ positions. The proportions of geo-entities are derived via semantic segmentation of satellite imagery and clustering of POI by categories.
  • Mixed-order Dependencies Representation. To model cross-region spatial and socioeconomic dependencies, we construct a heterogeneous graph connection that reflects the mixed-order relationships from spatial, environmental and societal perspectives.
  • Graph Construction for Geospace. We combine the basic features and cross-region dependencies using a comprehensive graph. The basic features are incorporated into the node vectors, while the mixed-order relations are expressed through the edges and their corresponding weights.
  • Graph-based Indicator Inference. We introduce two approaches to inference regional indicators on the graph: self-supervised graph learning for creating region embeddings for light-weight deployment and end-to-end training to directly inference on the graph for peak performance.
3.1 Space Feature Representation
To effectively encode spatial, environmental, and societal signals into feature embeddings, we extract multi-modal geo-contextual data including satellite imagery, POI data, and the location of the regions.
[…]
$E^{i}_{pos}=(lon^{i}-lon^{0},lat^{i}-lat^{0})\Delta*D^{-1}$ (1)
where $\Delta$ is the transform vector decided by the coordinate system. $D$ denotes the scale of the grids, and for $1\textrm{km}\times 1\textrm{km}$ grids, $D=1$. It is noteworthy that the 1 $1\text{ km}^{2}$ scale can be easily downsampled to other scales for improved compatibility.
Environmental Feature Embedding: We utilize satellite imagery to mine the environmental features and expect the satellite imagery encoding to be efficient and concise. To achieve this, we draw inspiration from the European Space Agency (ESA) with ESA WorldCover Dataset (Zanaga et al., 2022), elaborate in Appendix B, and design a segmentation-based process to encode the satellite images instead of the most commonly used visual encoders such as Resnet. Given a satellite image $S_{i}$ of a region of interest $R_{i}$, we conduct semantic segmentation to obtain a series of geo-entities $\{entity_{1},entity_{2},...,entity_{j}\}_{j=1}^{J}$ within that region, such as developed areas, grass, and trees based on ESA Worldcover, leveraging their area proportion as environmental feature embedding as below:
$E_{env}=\{p_{ent_{1}},p_{ent_{2}},...,p_{ent_{j}}\}_{j=1}^{J}$ (2)
where $p_{ent_{j}}=\frac{A_{j}}{A_{R}}$, and $A_{j}$ is the area occupied by $entity_{j}$, $A_{R}$ is the area of region $R_{i}$, $J$ is the number of environmental entity types. The detail discussion for opting for the segmentation-based method, along with extra validation tests, are elaborated in Appendix C.
$E_{soc}^{\prime}=\{p_{poi_{1}},p_{poi_{2}},...,p_{poi_{k}}\}_{k=1}^{K},\quad p_{\text{poi}_{k}}=\frac{|C_{k}\cap D_{POI}^{R_{i}}|}{|D_{POI}^{R_{i}}|}$
where $K$ is the number of POI categories, $C_{k}=\{poi_{j}\in D_{POI}\mid\text{category}(poi_{j})=k\}$. However, directly employing the proportion of different POI categories $E_{soc}^{\prime}$ as societal features can still be problematic since it cannot distinguish between areas with high POI density (usually with a higher degree of socialization) and those with relatively low POI density. Typically, a higher POI density region is more significantly influenced by its social characteristics. Thus, we convert $E_{soc}^{\prime}$ into the final societal feature embedding $E_{soc}$ by applying a Social impact factor $f$ as shown below:
$E_{soc}=f\cdot E_{soc}^{\prime},\quad f=log(D^{R_{i}}_{POI}+1)$ (3)
3.2 Mixed-order Dependencies Representation.
Region indicators are also significantly influenced by various cross-region dependencies. Our objective is to jointly model pairwise second-order spatial relationships between regions, as well as higher-order dependencies among groups of regions that exhibit similar environmental or societal characteristics, using a heterogeneous graph structure.
According to hypergraph theories (Liang et al., 2022; Feng et al., 2019), as elaborated in Appendix D.1, we employ a heterogeneous graph structure to represent cross-region relations. As shown in Figure 4, we utilize heterogeneous nodes in the graph structure as transfer nodes to facilitate high-order message passing (equivalent to hypergraph edges). Given a geospace $\mathcal{R}$, we construct a weighted heterogeneous graph $\mathcal{G}$, which contains three types of nodes: region nodes, environmental entity nodes, and societal entity nodes. Specifically, region nodes correspond to the grid cells of $\{r_{1},r_{2},\cdots,r_{N}\}$; environmental entity nodes represent the $J$ distinct geo-entity classes detected from satellite image $S_{i}$; and societal entity nodes comprise the $K$ different POI categories. The model structure can be found in Figure 3. We will elaborate on how second-order and high-order information is represented.
Second-Order Relation Representation: Second-order relations capture the pairwise spatial adjacency between regional nodes in $\mathcal{G}$. We construct undirected edges $\mathcal{E}_{RNR}$, Region Nearby Region (RNR) between regional nodes whose corresponding grid cells are spatially adjacent in a $3\times 3$ grid, thereby encoding these local second-order dependencies.
High-Order Relation Representation: Multi-view graphs (Geng et al., 2019; Zou et al., 2025) represent the relations between regions through additional edges and are confined to second-order regional connections. To fully capture the higher-order relations among urban regions, we explicitly model these higher-order interactions utilizing heterogeneous environmental entity nodes $\mathcal{V}_{env}$ and societal entity nodes $\mathcal{V}_{soc}$ within $\mathcal{G}$. Specifically, we construct heterogeneous weighted edges $\mathcal{E}_{ELR}$ and $\mathcal{E}_{SLR}$ through $\mathcal{V}_{env}$ and $\mathcal{V}_{soc}$ to associate each regional node with its constituting environmental entities (e.g., water, vegetation) and societal entities (e.g., educational, commercial POIs). For $p_{ent_{it}}\geq\theta_{Env}$ or $f\cdot p_{poi_{ik}}\geq\theta_{Soc}$:
where $p_{ent_{ij}}$ represents the area proportion of geo-entity $j$ at region $i$, $p_{poi_{ij}}$ denotes the quantity proportion of POI category $k$ at region $i$ and $f$ is the social impact factor elaborated in Equation 3. $\theta_{Env}$ and $\theta_{Soc}$ serve as hyperparameters designed to optimize the graph structure for conciseness and efficiency. This edge formulation allows encoding higher-order relationships between regions exhibiting similar environmental/societal attributes, even if they lack spatial proximity. By combining the second-order spatial adjacency edges with these higher-order hyperedge associations, $\mathcal{G}$ holistically represents the mixed-order relational patterns within the region.
3.3 Graph Construction for Geospace
After deriving the basic space features and the inter-regional dependencies encoded in the heterogeneous graph $\mathcal{G}$, we use a model-agnostic graph framework to integrate these information for a comprehensive space representation $E_{space}$. Specifically, this phase can be formulated as:
$E_{space}=\text{GNN}(\mathcal{G}[E_{pos},E_{env},E_{soc}])$ (6)
In the graph, each regional node $v$ in $\mathcal{G}$ is represented with an initial node feature corresponding to its regional feature $E_{pos},E_{env},E_{soc}$ elaborated in Sec 3.1. Note that for simplicity, instead of utilizing hypergraph neural networks (Feng et al., 2019; Yang et al., 2022; Xia et al., 2022), we model hyperedges with the form of complete subgraphs in $\mathcal{G}$. Therefore, we apply Heterogeneous Graph Neural Networks (HGNNs) over $\mathcal{G}$ to update the node representations by iteratively aggregating and transforming multi-hop neighborhood information:
where $\mathcal{N}(v)$ denotes the neighbors (spatial and higher-order) of node $v$ in $\mathcal{G}$. After $L$ layers of updates, the final node embedding $\mathbf{x}_{v}^{(L)}$ integrates cues from the basic features as well as mixed-order inter-region dependencies.
3.4 Graph-based Indicator Inference
To effectively address various practical demands, we present two inference strategies: self-supervised graph learning, which learns fine-grained, generalizable region representations within the space, enabling lightweight deployment and transferability to diverse downstream tasks; and end-to-end training, which allows for direct inference on the graph to achieve optimal performance.
Self-Supervised Graph Learning: Inspired by CLIP (Radford et al., 2021), we pre-train our model using a contrastive learning paradigm where we maximize the similarity between a region $\mathbf{R}_{i}$ and its corresponding positive region groups $\mathbf{C}_{i}$, while minimizing its similarity with other region groups in the batch. Specifically, we sample regional nodes adjacent to $\mathbf{R}_{i}$ as well as nodes exhibiting similar intra-regional features to construct $\mathbf{C}_{i}$. During training, we leverage a GNN model $\mathcal{M}_{pretrain}$ to obtain embeddings for $\mathbf{R}_{i}$ and all nodes in $\mathbf{C}_{i}$. The embeddings of the sampled positive regional nodes are further pooled to obtain the unified positive region representation $\mathbf{e}_{j}$. The process of generating embeddings for $\mathbf{R}_{i}$ and $\mathbf{C}_{i}$ is conducted as follows:
We further define a similarity function $f_{\mathbf{score}}$ to measure the similarity between the representations $\mathbf{e}_{i},\mathbf{e}_{j}$. $f_{\mathbf{score}}$ can be a simple dot product or a more complex metric and then employ an InfoNCE-based loss (Oord et al., 2018) to conduct contrastive learning:
where $f_{\mathbf{score}}(\mathbf{e}_{i},\mathbf{e}_{j})$ represents the score of positive pairs while $f_{\mathbf{score}}\left(\mathbf{e}_{i},\mathbf{e}_{n}\right)$ refers to the scores of negative pairs.
After obtaining the pre-trained model $\mathcal{M}_{pretrain}$ with pre-trained region embedding $\mathbf{E}_{pretrain}$ through self-supervised learning, we fine-tune it with a available samples by adding three trainable linear layers. The weights of these additional linear layers are updated during fine-tuning while the weights of $\mathcal{M}_{pretrain}$ and $\mathbf{E}_{pretrain}$ remain fixed:
$\hat{\mathbf{y}}=\text{MLP}(\mathbf{E}_{pretrain})$ (10)
End-to-End Training: For direct inference of socioeconomic indicators on the graph, the indicator inference can be regarded as a node label prediction task. In practice, we optimize all parameters from scratch using the task’s supervised signal. We adopt the HGNN model $\mathcal{M}$ to obtain embeddings for a given region $i$, which are then fed into a three-linear layer regression head denoted as $\mathbf{Regressor}$. All parameters are updated by minimizing the Mean Square Error (MSE) loss on the available sample nodes. The overall process can be formulated as follows:
$\hat{y}_{i}=\mathbf{Regressor}(\mathcal{M}(\mathbf{X}_{i})),\quad\mathcal{L}_{\text{MSE}}=\frac{1}{N}\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}$
4 Experiments
4.1 Experimental Setup
Datasets. The datasets used in this paper include satellite imagery, POI information and five region indicators (Population, GDP, Night Light, Carbon, $PM_{2.5}$) located in four representative cities in China: Beijing, Shanghai, Guangzhou, and Shenzhen. All indicators are provided at 1 $km^{2}$ scale by data servers. The satellite images and POI data are collected from Google Map API according to geospatial grids. The entity segmentation results for satellite imagery are collected from the European Space Agency. The overall dataset statistics can be listed in Table 1 and more information about region indicators can be found in Appendix E. We randomly mask 75% data as unlabeled samples for test, the available samples are divided to 80% for train and 20% for validation.
Table 1: Dataset statistics.
Dataset Coverage in Geospace Satellite Image POI Data
Bottom-left Top-right
Beijing 39.75°N, 116.03°E 40.15°N, 116.79°E 4,277 709,232
Shanghai 30.98°N, 121.10°E 31.51°N, 121.80°E 5,292 808,957
Guangzhou 22.94°N, 113.10°E 23.40°N, 113.68°E 8,540 805,997
Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 5,150 717,461
[…]
Methods GeoHG GeoHG-SSL UrbanVLP GeoStructural PG-SimCLR UrbanCLIP IDW UK
Metric $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE $\text{R}^{2}$ MAE RMSE
Beijing Carbon 0.894 0.112 0.211 0.887 0.161 0.224 0.717 0.363 0.467 0.705 0.398 0.482 0.402 0.651 0.764 0.604 0.558 0.598 0.734 0.367 0.489 0.745 0.355 0.454
Population 0.834 0.271 0.351 0.830 0.282 0.374 0.695 0.404 0.513 0.701 0.402 0.504 0.451 0.964 1.117 0.431 0.552 0.669 0.689 0.385 0.517 0.701 0.381 0.499
GDP 0.617 0.336 0.567 0.614 0.331 0.581 0.556 0.416 0.650 0.567 0.401 0.609 0.247 0.768 1.254 0.315 0.539 0.864 0.555 0.414 0.619 0.411 0.621 0.409
Night Light 0.861 0.239 0.311 0.850 0.244 0.353 0.531 0.394 0.629 0.488 0.429 0.657 0.369 0.404 0.728 0.420 0.457 0.700 0.668 0.304 0.414 0.681 0.301 0.384
$PM_{2.5}$ 0.921 0.064 0.160 0.920 0.065 0.169 0.641 0.484 0.954 0.694 0.306 0.482 0.398 0.624 0.845 0.533 0.556 0.699 0.791 0.139 0.271 0.809 0.129 0.234
Shanghai Carbon 0.915 0.157 0.290 0.912 0.162 0.312 0.716 0.392 0.529 0.688 0.413 0.557 0.298 0.712 0.914 0.671 0.426 0.569 0.696 0.395 0.563 0.713 0.388 0.532
Population 0.936 0.161 0.244 0.928 0.172 0.251 0.593 0.471 0.607 0.613 0.456 0.583 0.315 0.731 0.959 0.456 0.557 0.748 0.617 0.457 0.607 0.686 0.342 0.498
GDP 0.778 0.323 0.468 0.767 0.331 0.477 0.330 0.595 0.816 0.377 0.553 0.721 0.294 0.767 1.052 0.326 0.587 0.807 0.319 0.611 0.837 0.331 0.544 0.811
Night Light 0.898 0.222 0.311 0.891 0.234 0.347 0.457 0.494 0.667 0.442 0.517 0.685 0.308 0.566 0.768 0.387 0.511 0.709 0.444 0.571 0.756 0.464 0.517 0.729
$PM_{2.5}$ 0.866 0.120 0.379 0.836 0.150 0.394 0.486 0.497 0.654 0.527 0.398 0.592 0.303 0.617 0.895 0.444 0.518 0.774 0.592 0.342 0.554 0.643 0.291 0.542
Guangzhou Carbon 0.885 0.219 0.336 0.884 0.209 0.371 0.698 0.385 0.514 0.681 0.497 0.529 0.422 0.708 0.708 0.585 0.444 0.603 0.675 0.415 0.503 0.754 0.370 0.433
Population 0.871 0.244 0.274 0.855 0.255 0.299 0.665 0.441 0.481 0.687 0.433 0.485 0.303 0.954 0.972 0.533 0.567 0.687 0.674 0.471 0.483 0.695 0.433 0.438
GDP 0.715 0.371 0.532 0.712 0.366 0.574 0.436 0.541 0.764 0.439 0.533 0.699 0.282 0.897 1.264 0.440 0.546 0.762 0.411 0.625 0.748 0.476 0.611 0.763
Night Light 0.871 0.234 0.378 0.854 0.248 0.391 0.577 0.418 0.573 0.574 0.415 0.581 0.435 0.433 0.627 0.483 0.478 0.633 0.442 0.451 0.604 0.476 0.402 0.598
$PM_{2.5}$ 0.833 0.133 0.403 0.822 0.158 0.414 0.638 0.462 0.624 0.652 0.461 0.597 0.315 0.542 0.741 0.56 0.514 0.694 0.671 0.394 0.559 0.696 0.380 0.503
Shenzhen Carbon 0.926 0.128 0.290 0.912 0.162 0.304 0.659 0.418 0.568 0.647 0.431 0.581 0.257 0.683 0.816 0.562 0.483 0.571 0.241 0.577 0.726 0.189 0.634 0.887
Population 0.892 0.165 0.244 0.879 0.173 0.261 0.790 0.343 0.448 0.797 0.314 0.390 0.311 0.758 0.892 0.527 0.592 0.610 0.719 0.354 0.455 0.775 0.342 0.421
GDP 0.798 0.297 0.468 0.767 0.331 0.489 0.532 0.448 0.676 0.517 0.455 0.699 0.307 0.895 1.003 0.508 0.464 0.693 0.534 0.417 0.625 0.553 0.404 0.623
Night Light 0.942 0.149 0.245 0.939 0.155 0.268 0.457 0.459 0.667 0.445 0.488 0.682 0.454 0.358 0.588 0.387 0.511 0.709 0.413 0.513 0.719 0.396 0.528 0.758
$PM_{2.5}$ 0.906 0.116 0.298 0.905 0.117 0.331 0.566 0.494 0.598 0.597 0.451 0.491 0.323 0.613 0.883 0.430 0.586 0.762 0.653 0.398 0.426 0.749 0.345 0.391
Baselines. For comparison, we select two traditional interpolation methods (Inverse distance weighted (IDW) (Zimmerman et al., 1999; Lu and Wong, 2008) and Universal Kriging (UK) (Brus and Heuvelink, 2007)), the state-of-the-art LLM-enhanced methods ( UrbanCLIP (Yan et al., 2024)) and conventional socioeconomic indicator prediction methods (UrbanVLP (Hao et al., 2024), PG-SimCLR (Xi et al., 2022)) and graph-based approach (GeoStructual (Li et al., 2022)). Details about baselines is provided in Appendix F.
[…]
4.2 Comparison with State-of-the-Art Methods
We conduct an empirical assessment of various models across the four datasets, and the results of each model are detailed in Table 2. From these tables, we have two key findings: 1)Both GeoHG and GeoHG-SSL outperform all competing baselines over the 5 datasets for 4 cities. For instance, GeoHG surpasses the previous SOTA performance in Beijing, achieving $R_{2}$ improvements of +16.7%, +14.9%, +6.1%, +37%, +33% in Carbon, Population, GDP, Night Light and $PM_{2.5}$ tasks respectively. Consistent trends are observed across other cities, underscoring GeoHG’s stable accuracy. 2) Most socioeconomic indicator prediction methods perform even worse than traditional spatial interpolation methods on indicator inference tasks. GeoHG captures both spatial geometry and the underlying socioeconomic dynamics of geospace, resulting in lower data reliance compared to existing methods.
Figure 5: Visualization of Population Inference Results.
Besides, to better demonstrate the advantages of GeoHG’s non-continuous inference on large geospace, we evaluate the model against baseline methods on population inference in an expanded area of Shenzhen, using only 5% available data (863 points) to infer the population of remaining 16,401 $km^{2}$ space, as illustrated in Figure 5. The model accurately captures the true distribution, substantially outperforming traditional spatial interpolation methods like IDW and UK, which deviate significantly. This highlights that traditional methods, which rely solely on spatial relationships for interpolation, fail to model the intricate socio-environmental dependencies that are critical for characterizing the distribution of real-world socioeconomic indicators. In contrast, the effective joint modeling of geospace by GeoHG enables accurate indicator inference more than only spatial geometric dimensions.
Figure 6: Results of Ablation Study on $R^{2}$ Metric.
4.3 Ablation Study & Interpretation Analysis
Effects of Core Components. To examine the effectiveness of each core module, we conducted an ablation study based on the following variants: a) w/o Env, which excludes environmental information from the satellite imagery for embedding; b) w/o Soc, which omits POI entities for embedding; and c) w/o Pos, which does not utilize the corresponding location information of the regions. The $R^{2}$ results for five tasks in four cities are shown in Figure 6. We observe that removing location information significantly degrades performance in all tasks. Additionally, environmental features have a substantial impact on model accuracy for the PM2.5, Night Light, and GDP tasks. These results align with the intrinsic distribution characteristics of the task datasets. For instance, PM2.5 and night light data are primarily related to environmental elements such as vegetation and infrastructure. Conversely, data for carbon estimation is mainly derived from societal activity metrics. The analysis of the four cities reveals distinct distribution patterns across the five key indicators, highlighting the heterogeneity in geospatial characteristics. However, the robust performance of our model, which is constructed through the synthesis of three distinct perspectives, remains evident.
Effect of Graph Construction. Our framework utilizes heterogeneous graph and hyperedges between regions to reflect the geospatial high-order relations. To validate the effectiveness of high-order relation representation, we compare GeoHG against two variants: GeoHG-MLP which discards graph structure and relies only on intra-region feature representation, employing MLP for regression; GeoHG-Mono which keeps edges in adjacent regions while discarding hyperedges for high-order relations. The results presented in Figure 6 indicated that our GeoHG significantly benefits from efficient representation of complex mixed-order dependiencies in geospace, thereby resulting in enhanced performance.
Figure 7: Visualization of Geospatial Correlations learned by GeoHG. $(x,y)$ represents the numbering in the latitude and longitude.
Interpretation of Mixed-order Geospatial Correlations. To investigate the power of heterogeneous graph structure in capturing mixed-order dependencies for inference, inspired by GNNExplainer (Ying et al., 2019), we depict the trend of learned geospatial dependency for Carbon Emission inference in Guangzhou. By selecting the top $N$ important nodes by their weight in the GNN prediction process, as illustrated in Figure 7, we observe that GeoHG effectively captures the high-order correlations between the target region and remote region groups through message passing within the heterogeneous graph structure. The details about GNNExplainer and experiment results are introduced in Appendix F.2.
4.4 Qualitative Demonstration
To qualitatively assess the effectiveness of GeoHG in modeling geospace, we illustrate the regional dependency distribution by computing the cosine similarity obtained in GeoHG-SSL between different regions. As shown in Figure 8, we compare the relevant real-world situations of regions to examine our method’s ability to capture features and dependencies within the space. Due to page limitation, we introduce thr detailed explanation of the visualization results in Appendix G.
Figure 8: Regions’ similarity in GeoHG of Guangzhou. Darker indicate higher similarity with the region of interest.
4.5 Robustness Testing on Data Efficiency
To validate the robustness of GeoHG for inference with extremely limited data, we evaluate the model’s performance under masked ratios of 1%, 5%, 10%, and 15%, with results shown in Table 3. The results demonstrate the model’s strong data efficiency; with only 5% of the data, the model outperforms previous state-of-the-art socioeconomic indicator prediction models, such as UrbanVLP (Hao et al., 2024) and GeoStructural (Li et al., 2022), even though these models were trained on the entire training dataset across all tasks as described in their original papers.
Table 3: Data Efficiency of GeoHG in indicator inference. We take the average of $R^{2}$ results in 4 cities.
Masked Ratio Carbon Population GDP Light PM2.5
99% 0.720 0.713 0.535 0.703 0.694
95% 0.835 0.818 0.623 0.844 0.832
90% 0.850 0.846 0.635 0.864 0.856
80% 0.877 0.861 0.724 0.876 0.867